# Do computer vision foundation models learn the low-level characteristics of the human visual system?

### v1

@@ Abstract
Computer vision foundation models, such as DINO or OpenCLIP, are trained in a self-supervised manner on large image datasets. Analogously, substantial evidence suggests that the human visual system (HVS) is influenced by the statistical distribution of colors and patterns in the natural world, characteristics also present in the training data of foundation models. The question we address in this paper is whether foundation models trained on natural images mimic some of the low-level characteristics of the human visual system, such as contrast detection, contrast masking, and contrast constancy. Specifically, we designed a protocol comprising nine test types to evaluate the image encoders of 45 foundation and generative models. Our results indicate that some foundation models (e.g., DINO, DINOv2, and OpenCLIP), share some of the characteristics of human vision, but other models show little resemblance. Foundation models tend to show smaller sensitivity to low contrast and rather irregular responses to contrast across frequencies. The foundation models show the best agreement with human data in terms of contrast masking. Our findings suggest that human vision and computer vision may take both similar and different paths when learning to interpret images of the real world. Overall, while differences remain, foundation models trained on vision tasks start to align with low-level human vision, with DINOv2 showing the closest resemblance.
@@ 2 Related Work
Since the advent of deep learning, machine vision models based on foundation models [13, 30, 23] and DNNs [38, 28, 46] have successfully handled numerous advanced visual tasks. However, researchers have observed that machine vision operates differently from human vision. [16] revealed that standard convolutional neural networks (CNNs) trained on ImageNet are strongly biased toward texture recognition rather than shape, which contrasts with human visual patterns. [43] provided further evidence that deep neural networks (DNNs) differ significantly from the HVS, demonstrating poor robustness in object classification under 3D viewpoint changes and image distortions, and showing vulnerability to adversarial examples, which are rarely problematic for humans. [5] pointed out that DNNs performing well in benchmark tests share little overlap with biological vision mechanisms and fail to account for many findings in psychological studies of human vision. This highlights a clear distinction between machine and human vision, leading to the rise of interest in domain adaptation [7] and making networks robust to adversarial attacks [45].
@@ 3.1 Tested models and testing methodology
To test these models, we need to be able to compare pairs of images and assess the “perceived” difference between them. For example, for a pattern detection task, a pair of images could be a Gabor patch (test) and a uniform field (reference), as shown on the left of Figure 2. As such patterns are calibrated in physical light units in vision science, we generate these patterns as luminance maps, scaled in physical units of  cd/m2. These luminance maps are then mapped from the linear space to the sRGB color space using a display model (display peak luminance of 400 cd/m2) and fed into the image encoder of the foundation model for feature extraction. Note that the sRGB space is almost universally used to represent training datasets and is expected input for the tested encoders. To ensure that models can operate on small contrast values, we modified them to accept floating-point values (instead of 8-bit integers) as input. This was necessary as the quantization artifacts in 8-bit images are often larger than the detection thresholds of human vision.
@@
To compare the encoder responses with psychophysical data, we need to be able to map the image sampling frequency into the spatial frequency on the retina. For that, we select the effective resolution of 60 pixels-per-degree, which is typical for modern monitors. We note, however, that the choice of this parameter is arbitrary and the model similarity scores can be shifted by a small multiplier along the spatial frequency axis. The luminance maps are generated at the resolution of 224$\times$224 pixels, corresponding to the size of 3.7$\times$3.7 visual degrees.
For ColorVideoVDP, we do not use sRGB encoding as the metric can directly work on physical units. We directly use its quality score instead of $S_{\text{ac}}$. Note that the primary focus of this study is on foundation models, with the ColorVideoVDP metric used solely for comparison purposes.
@@
Figure 3: Gabors with different spatial frequencies (x-axis) and contrast (y-axis) used as the test images in the contrast detection tests (Section 4.1). “cpd” denotes cycles per degree. Note that the high-frequency patterns can be rendered with aliasing artifacts on the screen or in print — those were not present in our tests.
@@ 3.2 Model alignment score
As an example, we take the leftmost contour plot in row (a) of Figure 5. Each point on the contour plot corresponds to $S_{\text{ac}}$ between the test image as shown in Figure 3 and a uniform field of the same mean luminance. The dashed line represents human contrast detection data, predicted with castleCSF [3]. A well-aligned model should show one of the contour lines that follows the dashed castleCSF line. We cannot directly use the $S_{\text{ac}}$ values along the dashed line as the measure of alignment because some models result in $S_{\text{ac}}=0$ for most points near the detection threshold (they detect no difference). Therefore, instead, we rely on the measure of change in $S_{\text{ac}}$ in the neighborhood of the dashed line.
@@ 4 Experiments
Figure 4: Examples of test images for contrast detection and contrast masking. The detailed explanation of the stimuli can be found in the Supplementary. “Ach.” denotes achromatic.
Table 1: Key parameters for all our tests. Note that “Radius” does not apply when the pattern is not a Gabor.
@@
Spatial Frequency - Gabor RG 0.5 - 32 100 1 0.001 - 0.2
Spatial Frequency - Gabor YV 0.5 - 32 100 1 0.001 - 0.2
@@
Figure 5: Selected representative experimental results. Each row represents a test, and each column corresponds to a model, selected as the best-performing in their original tasks. (a)-(f): contour plots of contrast detection $S_{\text{ac}}$, with the ground truth castleCSF [3]. (g),(h): contour plots of contrast masking $S_{\text{ac}}$, with the ground truth from [14] and [15], respectively. (i): results from the contrast matching experiment, where different colors represent different $C_{r}$ values. The dashed lines are human results [18], while the solid lines are model-predicted results from Equation 2. Results for all models are in the supplementary material.
@@ 4.3 Supra-threshold contrast matching
$\operatorname*{argmin}_{c_{\text{t}}}\left(\frac{S_{\text{ac}}(F(\rho_{\text{r}},c_{\text{r}}),U)}{S_{\text{ac}}(F(\rho_{\text{r}},1),U)}-\frac{S_{\text{ac}}(F(\rho_{\text{t}},c_{\text{t}}),U)}{S_{\text{ac}}(F(\rho_{\text{t}},1),U)}\right)^{2}\,,$ (2)

### v2

@@ Abstract
Computer vision foundation models, such as DINO or OpenCLIP, are trained in a self-supervised manner on large image datasets. Analogously, substantial evidence suggests that the human visual system (HVS) is influenced by the statistical distribution of colors and patterns in the natural world, characteristics also present in the training data of foundation models. The question we address in this paper is whether foundation models trained on natural images mimic some of the low-level characteristics of the human visual system, such as contrast detection, contrast masking, and contrast constancy. Specifically, we designed a protocol comprising nine test types to evaluate the image encoders of 45 foundation and generative models. Our results indicate that some foundation models (e.g., DINO, DINOv2, and OpenCLIP), share some of the characteristics of human vision, but other models show little resemblance. Foundation models tend to show smaller sensitivity to low contrast and rather irregular responses to contrast across frequencies. The foundation models show the best agreement with human data in terms of contrast masking. Our findings suggest that human vision and computer vision may take both similar and different paths when learning to interpret images of the real world. Overall, while differences remain, foundation models trained on vision tasks start to align with low-level human vision, with DINOv2 showing the closest resemblance. Our code is available on https://github.com/caiyancheng/VFM_HVS_CVPR2025.
@@ 2 Related Work
Since the advent of deep learning, machine vision models based on foundation models [13, 31, 23] and DNNs [39, 28, 47, 29] have successfully handled numerous advanced visual tasks. However, researchers have observed that machine vision operates differently from human vision. [16] revealed that standard CNNs trained on ImageNet are strongly biased toward texture recognition rather than shape, which contrasts with human visual patterns. [44] provided further evidence that deep neural networks (DNNs) differ significantly from the HVS, demonstrating poor robustness in object classification under 3D viewpoint changes and image distortions, and showing vulnerability to adversarial examples, which are rarely problematic for humans. [5] pointed out that DNNs performing well in benchmark tests share little overlap with biological vision mechanisms and fail to account for many findings in psychological studies of human vision. This highlights a clear distinction between machine and human vision, leading to the rise of interest in domain adaptation [7] and making networks robust to adversarial attacks [46].
@@ 3.1 Tested models and testing methodology
To test these models, we need to compare pairs of images and assess the “perceived” difference between them. We adopt a methodology inspired by 2-alternative-forced-choice (2AFC) psychophysical experiments. For example, for a pattern detection task, a pair of images could be a Gabor patch (test) and a uniform field (reference), as shown on the left of Figure 2. As such patterns are calibrated in physical light units in vision science, we generate these patterns as luminance maps, scaled in physical units of  cd/m2. These luminance maps are then mapped from the linear space to the sRGB color space using a display model (display peak luminance of 400 cd/m2) and fed into the image encoder of the foundation model for feature extraction. Note that the sRGB space is almost universally used to represent training datasets and is expected input for the tested encoders. To ensure that models can operate on small contrast values, we modified them to accept floating-point values (instead of 8-bit integers) as input. This was necessary as the quantization artifacts in 8-bit images are often larger than the detection thresholds of human vision.
@@
To compare the encoder responses with psychophysical data, we need to be able to map the image sampling frequency into the spatial frequency on the retina. For that, we select the effective resolution of 60 pixels-per-degree, which is typical for modern monitors. We note, however, that the choice of this parameter is arbitrary and the model similarity scores can be shifted by a small multiplier along the spatial frequency axis. The luminance maps are generated at the resolution of 224$\times$224 pixels, corresponding to the size of 3.7$\times$3.7 visual degrees. It roughly aligned with the extend of the human foveal vision, and the stimuli span across multiple receptive fields of human vision and multiple patches/tokens of a foundation model.
For ColorVideoVDP, it can work directly on linear physical units, so conversion to sRGB was unnecessary. We directly use its quality score instead of $S_{\text{ac}}$. Note that the primary focus of this study is on foundation models, with the ColorVideoVDP metric used solely as a baseline for comparison purposes.
@@
Figure 3: Gabors with different spatial frequencies (x-axis) and contrast (y-axis) used as the test images in the contrast detection tests (Section 4.1). “cpd" denotes cycles per degree. Note that the high-frequency patterns can be rendered with aliasing artifacts on the screen or in print — those were not present in our tests.
@@ 3.2 Model alignment score
As an example, we take the leftmost contour plot in row (a) of Figure 5. Each point on the contour plot corresponds to $S_{\text{ac}}$ between the test image as shown in Figure 3 and a uniform field of the same mean luminance. The dashed line represents human contrast detection data, predicted with castleCSF [3]. A well-aligned model should show one of the contour lines that follows the dashed castleCSF line. We cannot directly use the $S_{\text{ac}}$ values along the dashed line as the measure of alignment because some models result in $S_{\text{ac}}=0$ for most points near the detection threshold (detect no difference). Therefore, instead, we rely on the measure of change in $S_{\text{ac}}$ in the neighborhood of the dashed line.
@@ 4 Experiments
Figure 4: Examples of test images for contrast detection and contrast masking. The detailed explanation of the stimuli can be found in the Supplementary. “Ach." denotes achromatic.
Table 1: Key parameters for all our tests. Note that “Radius" does not apply when the pattern is not a Gabor.
@@
Spatial Frequency - Gabor RG 0.5 - 32 100 1 0.001 - 0.12
Spatial Frequency - Gabor YV 0.5 - 32 100 1 0.001 - 0.8
@@
Figure 5: Selected representative experimental results. Each row represents a test, and each column corresponds to a model, selected as the best-performing in their original tasks. (a)-(f): contour plots of contrast detection $S_{\text{ac}}$, with the ground truth castleCSF [3]. (g),(h): contour plots of contrast masking $S_{\text{ac}}$, with the ground truth from [14] and [15], respectively. Different colored solid lines representing different $S_{\mathrm{ac}}$ values, where purple indicates the minimum difference ($S_{\mathrm{ac}}\to 0$). (i): results from the contrast matching experiment, where different colors represent different $C_{r}$ values. The dashed lines are human results [18], while the solid lines are model-predicted results from Equation 2. Results for all models are in the supplementary material.
@@ 4.3 Supra-threshold contrast matching
$\operatorname*{argmin}_{c_{\text{t}}}\left(S_{\text{ac}}(F(\rho_{\text{r}},c_{\text{r}}),U)-S_{\text{ac}}(F(\rho_{\text{t}},c_{\text{t}}),U)\right)^{2}\,,$ (2)